OpenStudy (anonymous):

15 medals (which ever subject you would like me to give you them in) Fan and Testimony! (DISCLAIMER: NOT A GRADED ASSINMENT!) Explain in at least 5 sentences how to find the circumference of a circle, a semicircle, and a quarter circle. How is finding circumference of a quarter-circle different than finding the perimeter of a quarter circle? Please write in complete sentences, don't just post the formulas. Thanks!

OpenStudy (anonymous):

@Donblue22 @sreekar369 @SuperflyTyGuy

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

I know its hard work but I just need help being "walked" step by step in it

OpenStudy (anonymous):

Circumference of a Quarter-Circle: \(C = \dfrac{2 \pi r}{4}\) Perimeter of a Quarter-Circle is the same thing as the circumference.. \(P = \dfrac{2 \pi r}{4}\) So there is no difference

OpenStudy (anonymous):

Is that writing in complete sentences? :D: P also medals on the way

OpenStudy (anonymous):

Yes, the circumference of a quarter-circle is just finding the circumference of a real circle and dividing by 4. And the Perimeter of the circle is found the same way.